In a certain group of 15 students, 5 have graphics calculators and 3 have a computer at home (one student has both). Two of the students drive themselves to college each day and neither of them has a graphics calculator nor a computer at home. A student is selected at random from the group.Find the probability that the student either drives to college or has a graphics calculator.
1
Expert's answer
2020-01-06T08:19:25-0500

A = set of students having computer , n(A) = 3

B = set of students having graphic calculator , n(B) = 5

C = set of students students who drive themselves to college each day , n(C) = 2


n(A⋂B) = 1

n(B or C)="n(B \bigcup C)" = 5 + 2 = 7

n(S) = 15

Required probability ="\dfrac{n(B⋃C)}{n(S)}" = "\dfrac{7}{15}"